PLEASE HELP ME SOLVE THIS EQUATION. 5X-2Y=-5, Y-5X=3.
THANK YOU SO MUCH. MRS. NEESE.

Answer by AnlytcPhil(1806)

PLEASE HELP ME SOLVE THIS EQUATION. 5X-2Y=-5, Y-5X=3.
THANK YOU SO MUCH. MRS. NEESE.

5X - 2Y = -5
Y - 5X = 3

Method 1: substitution:

                        5X - 2Y = -5
                         Y - 5X =  3

Solve the second equation for Y

Add +5X to both sides

                    Y - 5X = 3
                      + 5X    + 5X
                    --------------
                    Y      = 3 + 5x

Now replace Y in the first equation by (3 + 5X)

                   5X - 2Y = -5

            5X - 2(3 + 5X) = -5

Now solve it for X, First remove the parenthese: 
 
              5X - 6 - 10X = -5

                   -5X - 6 = -5
      
Add +6 to both sides

                       -5X = 1

Divide both sides by -5

                         X = -1/5

Now replace X by (-1/5) in the equation where
you solved for Y

                         Y = 3 + 5x
  
                         Y = 3 + 5(-1/5)

                         Y = 3 - 1

                         Y = 2

So the solution is  (X, Y) = (-1/5, 2)

----------------------------------------------
 
Method 2: Elimination (sometimes called "addition" or
"addition and subtraction")

                   5X - 2Y = -5
                    Y - 5X = 3

Rearrange the terms of the second equation so terms in the same letter
are lined up vertically:


                        5X - 2Y = -5
                       -5X +  Y =  3

Since the terms in X have coefficients which are equal in absolute
value but opposite in sign, they will cancel when added vertically:

                        5X - 2Y = -5
                       -5X +  Y =  3
                      --------------
                             -Y = -2
                              Y = 2

To cause the Y terms to cancel, leave the first equation as is but
multiply the second equation through by 2 so that they +Y term will
become +2Y and will cancel with the -2Y above it when adding the
two equations vertically:

                        5X - 2Y = -5
                     2[-5X +  Y =  3]

becomes:

                        5X - 2Y = -5
                      -10X + 2Y =  6
                     ---------------
                       -5X      =  1
                              X = -1/5

So the solution is  (X, Y) = (-1/5, 2) 
                       
Edwin
AnlytcPhil@aol.com